OpenStudy (barreraa):

Solve the given problem. Henry is four years older than half Linda’s age. If Linda is ten years old, how old is Henry? (Points : 1) 14 9 7 5

OpenStudy (hysusonic):

ok what do you think we have to do first in this equation?

OpenStudy (texaschic101):

H = (1/2)L + 4 L = 10 now we sub H = (1/2) 10 + 4 H = 5 + 4 H = 9

OpenStudy (barreraa):

is the answer 9

OpenStudy (texaschic101):

thats what I got

OpenStudy (barreraa):

The answer would be 9 because since linda is 10, half of ten is five then add four woulld be nine